OpenStudy (anonymous):

Find a simple formula for the nth term of the following sequence: 1, -2, 3, -4, 5, -6, ...

OpenStudy (anonymous):

make it alternate via \((-1)^n\) or \((-1)^{n+1}\) whichever is needed

OpenStudy (callisto):

1 = [(-1)^(1-1)](1) -2 = [(-1)^(2-1)] (2) 3 = [(-1)^(3-1)] (3) -4 = [(-1)^(4-1)] (4) 5 = [(-1)^(5-1)] (5) -6 = [(-1)^(6-1)] (6) . . . Two answers for this question :|

OpenStudy (anonymous):

it is simple 1,2,3,4,5,6,7,8......... but each even no is multiplied with minus so \[a_{n}=n(-1)(-1)^n\]

OpenStudy (anonymous):

but satellite doesn't work